i am a multiple of 5 and 10
i have 12 factors
the sum of my digit is 6
ring me

Answer by JBarnum(2146)

60 factores: 1-60, 2-30, 3-20, 4-15, 5-12, 6-10
meaning 1,2,3,4,5,6,10,12,15,20,60 thats 12 factors
60: 6+0=6
60/10=6
60/5=30
so the number your looking for has to be 60